OpenStudy (anonymous):

To calculate wavelength is it wavelength = frequency/velocity or wavelength = velocity/frequency?

OpenStudy (ujjwal):

wavelength=velocity/frequency

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

By the way, how do I calculate the length (wavelengths) of a tuning fork with f=512 Hz in a closed air column?

OpenStudy (anonymous):

since velocity=distance/time and frequency=1/time so for light; velocity=wavelength x frequency and then wavelength=velocity/frequency

OpenStudy (anonymous):

ohh.. i see, thank you!

OpenStudy (ujjwal):

any clue for the value of velocity? @milliex51

OpenStudy (anonymous):

I'm also confused with how my first resonant point from my investigation was 17 cm in length and I'm supposed to find the length (wavelenghts) with a tuning for of 512 Hz..

OpenStudy (anonymous):

The velocity is 346. 16: v = 332 m/s + (0.59 m/s / degrees C) (24 degrees Celsius)

OpenStudy (anonymous):

*346.16 m/s

OpenStudy (ujjwal):

and \[V=f \lambda\] so carry on..

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

one way to think about this is to consider the units. Wavelength is a distance (meters). Velocity is meters/second. Frequency is cycles per second. So wavelength = velocity / frequency = m/s divided by 1/s = meters

OpenStudy (anonymous):

Speed of light = C = Frequency x Wavelength. So C/F = W or C/W=F